OpenStudy (anonymous):

Solve using addition method. 2x-y=7 x+2y=1

OpenStudy (anonymous):

Multiply the first eqn by 2 throughout and add the two eqn to eliminate the y term

OpenStudy (anonymous):

2x-y=7 therefore 4x-2y=14....(1) x+2y=1.....(2) add 1 and 2 you will get 5x=15 as 2y will be canceled out. so x=3. putting x in eqn 2 you will get 2y=-4=> y=-2.

OpenStudy (anonymous):

4x -2y = 7 x + 2y = 1 Adding both 5x = 8 So x = 8/5

OpenStudy (anonymous):

Sorry, it is 4x -2y = 14 x + 2y = 1 Adding both 5x = 15 So x = 15/5 ie 3

OpenStudy (anonymous):

Then as dipankar has done